The width of a rectangular swimming is pool is 79 ft. If the perimeter is 324 ft, what is the area of the pool?

Answer by palanisamy(496)

Let the length of the swimming pool = x
Given, its width = 79 ft
Perimeter is 2*(x+79) = 324
x+79 = 324/2
x = 162-79
x = 83 ft
Area = 79*82 = 6478 sq ft